OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)) I don't know the answer to this problem

OpenStudy (anonymous):

f(g(x))=f(-x+3)=9*(-x+3)+3 do you know the rest? :)

OpenStudy (anonymous):

that's all it says

OpenStudy (anonymous):

no, what i meant to say is that only thing left is to multiply and sum after that :) do you know how to do that by yourself?

OpenStudy (anonymous):

so will the answer be -9x+27+3=-9x+30??

OpenStudy (anonymous):

yes :) well done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're wellcome

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). so does that mean -40x + 24 would be the answer for this problem

zepdrix (zepdrix):

Woops, I think you got your functions a little mixed up Stephan. f = 9x-2 So when you plug g into f, = 9g-2 = 9(-x+3)-2 Should be -2 on the far right, not +3. Small typo probably :)

OpenStudy (anonymous):

oh yes, i mixed it up, sorry :)